Parth (parthkohli):

OpenStudy in real life problem: A user has 9 letters and then 2 numbers in his/her username. If you know that the first letter is 's' and the last number is '3', then how many combinations are possible for the name? Assume that all letters are small.

Parth (parthkohli):

Hint: the first and last are fixed values.

Parth (parthkohli):

The solution would be posted in a while.

OpenStudy (slaaibak):

9 * 26^8 * 10

Parth (parthkohli):

How did you get the answer? explain my friends

OpenStudy (maheshmeghwal9):

satellite73 but don't know combination

OpenStudy (aravindg):

26^8 *10

OpenStudy (aravindg):

@ParthKohli why did u post the answer so fast ??

OpenStudy (slaaibak):

You are forgetting there are 9 possible arrangements.

Parth (parthkohli):

lol but satellite73 is correct =))

OpenStudy (aravindg):

i waas going to xplain?

Parth (parthkohli):

9 possible rearrangements?

Parth (parthkohli):

Dw lol this was an easy problem.. I'll post a much more difficult one after this

OpenStudy (aravindg):

@ParthKohli bt pls post the soln only after sufficient time

Parth (parthkohli):

I have deleted it.

OpenStudy (kropot72):

25P8 * 10 = 25 * 24 * 23 * 22 * 21 * 20 * 19 * 18 * 10

OpenStudy (aravindg):

nah ..nw evryone have read it :(

Parth (parthkohli):

No wait... letters can repeat.. I didn't say they can't

OpenStudy (slaaibak):

# x x x x x x x x is not the same as x # x x x x x x x

Parth (parthkohli):

And then 2 numbers. I guess I framed the problem wrong.

Parth (parthkohli):

It's like s x x x x x x x x # 3

Parth (parthkohli):

I had taken satellite as an example, I'm sorry I framed it wrong though

OpenStudy (slaaibak):

oh... yeah you should state clearly that the numbers appear at the end, that was a bit ambiguous.. if the numbers should be in the last two spots, its 26^8 * 10, yes.

Parth (parthkohli):

Sorry for that sir.

OpenStudy (callisto):

number of English alphbet = 26 number of number = 10 The first and the last one is fixed that is s _ _ _ _ _ _ _ 3 Where _ can be a letter or a number Case one: the number is the second one, that is s (number) (letter)x6 3 Number of combinations = 1 times 10 times 26^6 times 1 = 10 x 26^6 Case two, the number is the third one, that is s (letter) (number) (letter)x5 3 = 1 times 26 times 10 times 26^5 times1 = 10 x 26^6 ... Case seven, the number is the 8th one, that is s (letter)x6 (number) 3 = 1 times 26^6 times 10 times 1 = 26^6 times 1 Add all the cases, P required = 7 x 10 x 26^6 This must be wrong... I hardly solve a probability/combination question correctly

Parth (parthkohli):

@Callisto can you give a medal to slaibak?

OpenStudy (aravindg):

i will :)

Parth (parthkohli):

Thanks :)

OpenStudy (callisto):

You can give the medal to slaibak :)

OpenStudy (aravindg):

ya @Callisto is going to give me :P